how do I find the area of a equilateral triangle by only knowing one side of it which is ten ft. I know this is not where I submit this problem but I have a hard time finding the right place to go thank you
Answer by ankor@dixie-net.com(22740)

how do I find the area of a equilateral triangle by only knowing one side of it which is ten ft
:
Draw the triangle, draw the height of the triangle which you need to find the area.
:
Notice that it bisects the base giving you 2 right triangles with one side of 5 ft and the hypotenuse of 10. The third side is the height of the triangle, we can find that by using the pythagorean theorem:
:
Let the unknown side be = h
5^2 + h^2 = 10^2
25 + h^2 = 100
h^2 = 100 - 25
h^2 = 75
h = SqRt(75)
h = 8.66 ft
:
Find the area:
A = .5*b*h
A = .5*10*8.66
A = 43.3 sq ft